Suppose f(0)=2, f’(0)=7, f(1)=5, f’(1)=8, g(1)=1, g’(1)=-9 and H(x)=f(ln(x))+ln(g(x)). Find dH/dx, x=1
What is H’(1)?
1
Expert's answer
2019-10-10T13:52:00-0400

Q96195


Solution:

"d\/dx(f(x)+g(x))=f'(x)+g'(x)" -------(1)

And for a composite function f(g(x)), its's derivative is defined as:

"df\/dx=f'(g(x))*g'(x)" -----(2)

Moreover, "d\/dx(ln(x))=1\/x" -----(3)

Also, given:

"g(1)=1" ----(4)

"g'(1)= -9" ----(5)

"f'(0)=7" ----(6)


Now, "H(x)=f(ln(x))+ln(g(x))"

thus, using (1), (2) and (3),

"H'(x)=" "dH\/dx=f'(ln(x))\/x +g'(x)\/g(x)"

hence, "H'(1) = f'(ln(1))\/1 +g'(1)\/g(1)" (since, "ln(1) = 0" )

"=f'(0) + g'(1)\/g(1)"

"= 7 + (-9\/1)" (using (4), (5) and (6))

"= 7-9"

= -2 (Answer)